You normally drive on the freeway at an average speed of 105 𝐾𝑃𝐻 between San Diego and Los Angeles that takes 2β„Ž and 20 π‘šπ‘–π‘›. However, on a Friday afternoon, heavy traffic slows you down, and you drive the same distance at an average speed of only 70 KPH. How much longer does the trip take?


1
Expert's answer
2021-09-25T13:11:01-0400
"\Delta t=t_2-t_1\\\Delta t=\frac{105}{70}\frac{120+20}{3}-\frac{120+20}{3}=23\ min"